NOAA Office of Satellite and Product Operations


This NOAA site will no longer provide GOES-East imagery. For access to high resolution GOES-East imagery from GOES-16, please go to the site: We apologize for the inconvenience and appreciate your patience.

MetOp-A HIRS Subsystem Summary

Component Status:

CHANNEL 01Spectral Channel #1, 14.95 micro metersGreen
CHANNEL 02Spectral Channel #2, 14.71 micro metersGreen
CHANNEL 03Spectral Channel #3, 14.49 micro metersGreen
CHANNEL 04Spectral Channel #4, 14.22 micro metersGreen
CHANNEL 05Spectral Channel #5, 13.97 micro metersGreen
CHANNEL 06Spectral Channel #6, 13.64 micro metersGreen
CHANNEL 07Spectral Channel #7, 13.35 micro metersGreen
CHANNEL 08Spectral Channel #8, 11.11 micro metersGreen
CHANNEL 09Spectral Channel #9, 9.71 micro metersGreen
CHANNEL 10Spectral Channel #10, 12.47 micro metersGreen
CHANNEL 11Spectral Channel #11, 7.33 micro metersGreen
CHANNEL 12Spectral Channel #12, 6.52 micro metersGreen
CHANNEL 13Spectral Channel #13, 4.57 micro metersGreen
CHANNEL 14Spectral Channel #14, 4.52 micro metersGreen
CHANNEL 15Spectral Channel #15, 4.47 micro metersGreen
CHANNEL 16Spectral Channel #16, 4.45 micro metersGreen
CHANNEL 17Spectral Channel #17, 4.13 micro metersGreen
CHANNEL 18Spectral Channel #18, 4.00 micro metersGreen
CHANNEL 19Spectral Channel #19, 3.76 micro metersGreen
CHANNEL 20Spectral Channel #20, 0.69 micro metersGreen
ELECTRONICSInstrument ElectronicsGreen
SCAN MOTORInstrument Scan MotorGreen

Subsystem History:

Date Time Component Description
10/23/201115:55:52 Recovery from PLSOL
10/22/201122:38:23 PLSOL Occurred
09/29/201103:50:00ELECTRONICSHIRSReactivation. Data outage: 28/09/11 07:24 to 29/09/11 03:50.
09/29/201103:50:00 HIRSReactivation. Data outage: 28/09/11 07:24 to 29/09/11 03:50.
09/28/201107:24:00ELECTRONICSHIRS parked at internal warm target to support maneuvers.
09/28/201107:24:00 HIRS parked at internal warm target to support maneuvers.
10/07/201005:45:56 HIRSReactivation. Data outage: 2010-278-07:31:46 to 2010-279-05:45:56
10/06/201007:31:46 HIRS parked at internal warm target to support maneuvers.
09/17/200914:50:33 HIRS Internal Warm Target Reactivation
09/17/200909:13:44 HIRS Internal Warm Target Deactivation
04/09/200806:03:33 OOP Maneuver Completion: HIRS commanded Position Disable, returned to Full Scan mode. Science Packets are nominal.
04/08/200807:35:06 OOP Maneuver Configuration: HIRS Scanner to Internal Warm Target (IWT) Position. Calibration Mode Enabled. Science Packets are filled with calibration data.
03/21/200814:25:00 PLSOL Recovery
03/19/200823:21:13 PLSOL occurred
01/17/200821:17:00 PLSOL Recovery
01/16/200813:46:00 PLSOL occured
09/19/200714:30:59 Turn on HIRS after autonomous turn off
09/17/200706:05:00 Data interruption due to autonomous payload switch off. Investigation and recovery underway.
05/22/200712:00:00CHANNEL 18MetOp-2 went operational
05/22/200712:00:00CHANNEL 01MetOp-2 went operational
05/22/200712:00:00ELECTRONICSMetOp-2 went operational
05/22/200712:00:00CHANNEL 10MetOp-2 went operational
05/22/200712:00:00CHANNEL 11MetOp-2 went operational
05/22/200712:00:00CHANNEL 12MetOp-2 went operational
05/22/200712:00:00CHANNEL 13MetOp-2 went operational
05/22/200712:00:00CHANNEL 14MetOp-2 went operational
05/22/200712:00:00CHANNEL 15MetOp-2 went operational
05/22/200712:00:00CHANNEL 16MetOp-2 went operational
05/22/200712:00:00CHANNEL 17MetOp-2 went operational
05/22/200712:00:00CHANNEL 19MetOp-2 went operational
05/22/200712:00:00CHANNEL 02MetOp-2 went operational
05/22/200712:00:00CHANNEL 20MetOp-2 went operational
05/22/200712:00:00CHANNEL 06MetOp-2 went operational
05/22/200712:00:00CHANNEL 07MetOp-2 went operational
05/22/200712:00:00CHANNEL 09MetOp-2 went operational
05/22/200712:00:00SCAN MOTORMetOp-2 went operational
05/22/200712:00:00CHANNEL 08MetOp-2 went operational
05/22/200700:00:00CHANNEL 03MetOp-2 went operational
05/22/200700:00:00CHANNEL 04MetOp-2 went operational
05/22/200700:00:00CHANNEL 05MetOp-2 went operational
05/22/200700:00:00 MetOp-2 went operational